• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 16:05:27 on July 26, 1999

    Entry number 23739

    keyword=Beam Time Accounting
    Hall A

    7/26/1999 0800-1600
    Experiment #: E91-010
    Shift Leader: B. Wojtsekhowski
    Shift Workers: K. Aniol, D. Armstrong, R. Michaels

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    1.0 1.0 47 0.0 0.0 0.0 5 5 2.95
    Beam Available
    but Not in Use
    0 0 0 0 0 0 0 0 0
    Beam Not Available
    or Unacceptable
    0 0 13 1.0 1.0 1.0 55 55 5.05
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 8
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 0 0
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: 8:00 - RCS tests (no spectrometer of HAPPEX data) 9:30 - go to Pb target tests; vary rate and watch temperature beam off at 11:00; problems with 45 MeV dipole

    Data Entry by: D. Armstrong